OpenStudy (anonymous):

write in point slope form the equation of the line that passes through the given point and has the given slope (5,2) m= -2

sam (.sam.):

y-y1=m(x-x1)

sam (.sam.):

y-2=-2(x-5) y=-2x+10+2 y=-2x+12